OpenStudy (anonymous):

find the xintercepts and vertext of the porabola y=x^2+2x-24

OpenStudy (anonymous):

y+25=(x+1)^2 ===> vertex =(-1,-25) , x-intercept is x=4 or x=-6

OpenStudy (anonymous):

thank yoU!